You are here: Foswiki>PartnerObserving Web>TechNightLogs>December21UT (21 Jul 2009, NormCushing)Edit Attach
21 December 2008 UT

PartnerObserving Log

morning & afternoon:
  • Fernando & Vincenzo noticed this morning that the Blue Science controller fans, which they found not working but had revived yesterday afternoon (not replaced as I had written earlier) had slowed down. Concerned they would not last, they replaced these with the fans from the Red Spare controller.

  • Andrea made changes to the LBC software

17:00: Turning the LBC on and testing.

17:30 Mostly clear skies - only some cirrus. Vent doors opened and Aaron is opening the enclosure.

20:00 Slewed to SA92 to start collimation and focus. Observed standard

20:35 Collimate again - note LBCFPIA is predicting seeing 1" (B) and 0.8" (R)

20:40 Check focal plane tilt using RB_rVsuperfoc.ob and RB_rVsuperfoc180.ob. PA=0 set: 033959-034225 gives best focus for chip 2 at:
 at top:      0.103 fwhm = 3.35
 at center:  0.117 fwhm=3.41
 at bottom: 0.138 fwhm = 3.78
PA=180 set:
 at top:      0.075 fwhm = 3.89
 at center: 0.0994 fwhm=3.95 
 at bottom: 0.122  fwhm = 4.27
So, best focus at the bottom of the chip 2 is larger (30-47 microns) than at the top, regardless of rotator angle. This indicates a tilt - I am not sure how this compares to the previous mounting of the red camera. I will check the blue camera data next.

Analyzing chips 1, 3 and 4 for PA=0:
Chip1top:      0.117 4.00
chip1 center  0.140 4.36
chip 1 bottom 0..160 4.99

Chip3 top: 0.110 3.89
chip3 cen: 0.115 3.69
chip3 bot:  0.133 4.01

bottom to top difference in best-focus is 27 microns for PA=0 and -27 mic for PA=180.

21:30 Numerous times the right main mirror failed to focus - it reaches position but issues this message just before. adding later maybe it does not reach position - Aaron has had to click 'collimate' to make the mirror reach position.

22:03 1.08" in Red and 0.908" in Blue.

22:30 Taking data to measure filter focus offset - elev = 60 deg. Start with rV and took 2 sets, because we had many right main mirror failed to focus errors during the first set. First set gives best focus = 0.078 fwhm=2.74

22:50 Take 5s image pair to check rotator center: Take a 5 s image with PA=0 rV 055053 (B) and 055220 (R). Right mirror problems We collimated just before taking this pair and the B image was finished before the red even started. We got a right main mirror failed to focus error. While the values on PSF Right were still in yellow, the Red exposure was taken. The images have signficant coma. We took a 2nd pair of images with PA=180, 055413 (B) and 055411 (R), the Red images are round. So it is does seem that the mirror really did not get into position before 055220 was started.

23:30 On phone with John to troubleshoot red mirror 'stickiness'. Vincenzo continuing to take data.

23:57 Using superfoc at 0 and 180 PA, 4th image in each which closest to best focus (034149 for pa=0 and 034746 for pa=180), to estimate rotator center. Had to imcopy[LBCCHIP2] to new image. rotcenter result is below:
      Old rotator center: 1079. 2893.
      New rotator center:
      ##################################################
      Xc=1071.96 Yc=2916.03 from geomap.
      Xc=1071.95 Yc=2916.12 from averages.
      ##################################################
      Images are 2304 pixels wide

00:05 Back to observing. Aaron is filing issue traks - both about the DX mirror collimation and actuator #115 which complains about firmware. We'll have to keep an eye on the mirror position and when the exposure starts.

02:00 The DX mirror has just not moving to position - many error messages. See issue #1912.

03:40 Polaris monitor reading 0.9" --- seeing is very good.

04:00 An OB which has 30 min open shutter time is taking over 1 hr to complete because of the red mirror 'stickiness' --- images B 081927-090827 are each 6min images at Us and the R is taking r and i images at each dither. Going to shut off red camera.

06:15 notice 6/5 dithers. Andrea has implemented a way to recover from a failure and then to check for images which are lost through the error. If there are images lost, then the OB restarts and skips dither positions with good data. The OB Progress in this case, if, e.g. the number of dithers is 5 will report, 6/5.

06:20 Starting to take flats. VV 133621 and 133730 have ~50K counts. Note that the ds9 display from file->open is very slow. It takes 45 seconds to open the browse box (network problem or other cause?) This slowness is noted on both lbtmu05 as okuhn and on lbtmu03 as lbcobs. top on lbtmu05 show;
Cpu(s):  2.6% us,  1.0% sy,  0.0% ni, 33.1% id, 63.2% wa,  0.0% hi,  0.0% si
Mem:   1035148k total,  1022448k used,    12700k free,     1508k buffers
Swap:  2621432k total,    16372k used,  2605060k free,   680812k cached

06:43 Aborted 30s VV flat sequence and changed exposure time to 10s. Saturated at center.

06:53 Starting 1s Us flats.

-- OlgaKuhn - 21 Dec 2008
Topic revision: r6 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback